OpenStudy (anonymous):

:EXPONENTIAL FUNCTIONS: Provide the solution to the equation 25 = square root of 5 (2x + 6)

OpenStudy (anonymous):

\[25=\sqrt{5} ^{2x+6}\]

OpenStudy (anonymous):

\[25=\sqrt{5}^4\] is a start

OpenStudy (anonymous):

alright so \[5^{2} = 5^{x +3}\]

OpenStudy (anonymous):

then if \[\sqrt{5}^4=\sqrt{5}^{2x+6}\] you know \[4=2x+6\] and you can solve for \(x\)

OpenStudy (anonymous):

@mebs okay so I get that then ?

OpenStudy (anonymous):

well than its just a matter of exponents because the base is the same... 2 = x + 3 can you do it

OpenStudy (anonymous):

just making sure you took out the 2 from 2x+6 and then made the other side 5^2 by ?

OpenStudy (anonymous):

Satellite also has the correct answer as well... yes that is correct

OpenStudy (anonymous):

Here is what my though process was.... 25 = 5^2 that's simple right.... squareroot5 = 5^(1/2) multiply 1/2 by (2x +6) distribute the 1/2 so you get x +3 alright... Now its the same base 5^2 = 5^x+3

OpenStudy (anonymous):

THANKS!

OpenStudy (anonymous):

yw =))